# homogeneous_khovanskii.bib

### v1

@@ 3.3. Computing homogeneous MUVAK bases
Let $h\in\alpha^{-1}(f)$. If $\operatorname{mindeg}_{i}(h)=v_{i}(f)$ for all $i\in\{1,\dots,m\}$, then $f$ is faithfully representable in $f_{1},\dots,f_{k}$ via $h$. So, assume there is $i\in\{1,\dots,m\}$ with $\operatorname{mindeg}_{i}(h)<v_{i}(f)$, hence $\operatorname{in}_{i}(h)^{\mathrm{hom}}\in I^{\mathrm{hom}}+\langle t_{i}\rangle$ by assumption. Then there is $h^{\prime}\in\langle t_{i}\rangle$ with $\operatorname{in}_{i}(h)^{\mathrm{hom}}-h^{\prime}\in I^{\mathrm{hom}}$, so $\operatorname{in}_{i}(h)-h^{\prime\mathrm{deh}}\in I$. We may assume $\deg_{j}(h^{\prime})=\deg_{j}(\operatorname{in}_{i}(h)^{\mathrm{hom}})$ for all $j\in\{1,\dots,m\}$, so $\operatorname{mindeg}_{j}(h^{\prime\mathrm{deh}})\geq\operatorname{mindeg}_{j}(\operatorname{in}_{i}(h))$ and $\operatorname{mindeg}_{i}(h^{\prime\mathrm{deh}})>\deg_{i}(\operatorname{in}_{i}(h))$ as $t_{i}\mid h^{\prime}$. We now replace $h$ by $h-\operatorname{in}_{i}(h)+h^{\prime\mathrm{deh}}$. By continuing to increase the $\deg_{i}$/̄degree in the described way, we arrive at a polynomial $\tilde{h}\in\alpha^{-1}(f)$ with $\operatorname{mindeg}_{i}(\tilde{h})=v_{i}(f)$ and $\operatorname{mindeg}_{j}(\tilde{h})\geq\operatorname{mindeg}_{j}(h)$ for all $j\neq i$ after finitely many steps. We may now continue this process with another index $i^{\prime}$ for which $\operatorname{mindeg}_{i^{\prime}}(\tilde{h})<v_{i^{\prime}}(f)$ and eventually we obtain a polynomial $h^{\prime}\in R$ with $\alpha(h^{\prime})=f$ and $\operatorname{mindeg}_{j}(h^{\prime})=v_{j}(f)$ for all $j\in\{1,\dots,m\}$, so $f$ is faithfully representable. ∎
@@
Let $f_{1},\dots,f_{k}\in A$ be $\Delta$/̄homogeneous generators and let $I\coloneqq\ker(\alpha)$ and $J_{i}\coloneqq\ker(\beta_{i})$, $1\leq i\leq m$, with the morphisms $\alpha$ and $\beta_{i}$ as above. The set $\{f_{1},\dots,f_{k}\}$ is a $\Delta$/̄homogeneous MUVAK basis of $A$ with respect to $v_{1},\dots,v_{m}$ if and only if we have $J_{i}^{\mathrm{hom}}\subseteq I^{\mathrm{hom}}+\langle t_{i}\rangle$ for all $i\in\{1,\dots,m\}$.
@@
Assume that $\{f_{1},\dots,f_{k}\}$ is a $\Delta$/̄homogeneous MUVAK basis of $A$ with respect to $v_{1},\dots,v_{m}$. Fix $i\in\{1,\dots,m\}$ and let $h\in J_{i}$. As $J_{i}$ is a $\Delta$/̄homogeneous ideal, we may assume that $h$ is $\Delta$/̄homogeneous. Set $f\coloneqq h(f_{1},\dots,f_{k})$. By assumption, the polynomial $f$ is faithfully representable, so by Lemma 3.10, we have $h^{\mathrm{hom}}\in I^{\mathrm{hom}}+\langle\vartheta\rangle$ with $\vartheta=\prod_{j=1}^{m}t_{j}^{v_{j}(f)-\deg_{j}(h^{\mathrm{hom}})}$. By Lemma 2.15, we have $\operatorname{mindeg}_{i}(h)<v_{i}(f)$, so $t_{i}\mid\vartheta$. Hence $h^{\mathrm{hom}}\in I^{\mathrm{hom}}+\langle t_{i}\rangle$.
Assume now $J_{i}^{\mathrm{hom}}\subseteq I^{\mathrm{hom}}+\langle t_{i}\rangle$ for all $i\in\{1,\dots,m\}$. Let $f\in A$ be $\Delta$/̄homogeneous and pick $h\in\alpha^{-1}(f)$. If $\operatorname{mindeg}_{i}(h)\geq v_{i}(f)$ for all $i\in\{1,\dots,m\}$, then $f$ is faithfully representable, so assume that there is $i\in\{1,\dots,m\}$ with $\operatorname{mindeg}_{i}(h)<v_{i}(f)$. Then $\operatorname{in}_{i}(h)\in J_{i}$ by Lemma 2.15, so $\operatorname{in}_{i}(h)^{\mathrm{hom}}\in J_{i}^{\mathrm{hom}}\subseteq I^{\mathrm{hom}}+\langle t_{i}\rangle$. Hence $f$ is faithfully representable in $f_{1},\dots,f_{k}$ with respect to $v_{1},\dots,v_{m}$ by Lemma 3.12. ∎
@@
Input : $\Delta$/̄homogeneous generators $f_{1},\dots,f_{k}\in A\leq B$, valuations $v_{1},\dots,v_{m}:B\setminus\{0\}\to\mathbb{Z}$
Output : A $\Delta$/̄homogeneous MUVAK basis of $A$ with respect to $v_{1},\dots,v_{m}$
1 Initialize $\mathcal{B}\coloneqq\{f_{1},\dots,f_{k}\}$
@@ 2 Compute the ideals $I^{\mathrm{hom}}$ and $J_{i}^{\mathrm{hom}}$ for all $i\in\{1,\dots,m\}$
5           Write $J_{i}^{\mathrm{hom}}=\langle h_{1},\dots,h_{r}\rangle$ and =$\mathcal{B}=\{f_{1},\dots,f_{l}\}$
6           $\mathcal{B}\coloneqq\mathcal{B}\cup\{h_{j}^{\mathrm{deh}}(f_{1},\dots,f_{l})\mid h_{j}\notin I^{\mathrm{hom}}+\langle t_{i}\rangle,\leavevmode\nobreak\ j=1,\dots,r\}$
@@ 7           Update $I^{\mathrm{hom}}$ and $J_{i}^{\mathrm{hom}}$
We give an efficient method to compute the homogenization of an ideal with respect to the in general only non/negative weights $v_{i}(f_{j})$ in Appendix A.
@@
Let $\{\tilde{f}_{1},\dots,\tilde{f}_{s}\}\subseteq A$ be a $\Delta$/̄homogeneous MUVAK basis of $A$ with respect to $v_{1},\dots,v_{m}$ and let $f_{1},\dots,f_{k}\in A$ be $\Delta$/̄homogeneous generators. If $\tilde{f}_{i}$ is faithfully representable in $f_{1},\dots,f_{k}$ for all $i\in\{1,\dots,s\}$, then $\{f_{1},\dots,f_{k}\}$ is a $\Delta$/̄homogeneous MUVAK basis for $A$ with respect to $v_{1},\dots,v_{m}$.
@@
Let $f\in A$ be $\Delta$/̄homogeneous. Then $f$ is faithfully representable in $\tilde{f}_{1},\dots,\tilde{f}_{s}$, so there is $h\in K[Y_{1},\dots,Y_{s}]$ with $h(\tilde{f}_{1},\dots,\tilde{f}_{s})=f$ and $\operatorname{mindeg}_{i}(h)=v_{i}(f)$ for all $i$. Put $\tilde{h}\coloneqq h(h_{1},\dots,h_{s})\in K[X_{1},\dots,X_{k}]$, so $\tilde{h}(f_{1},\dots,f_{k})=f$. By definition, we have $\deg_{i}(Y_{j})=v_{i}(f_{j})$ for all $i=1,\dots,m$ and $j=1,\dots,s$. Hence $\operatorname{mindeg}_{i}(\tilde{h})\geq v_{i}(f)$ for all $i$ as required. ∎
@@
Algorithm 3 terminates after finitely many steps if and only if there exists a finite $\Delta$/̄homogeneous MUVAK basis of $A$ with respect to $v_{1},\dots,v_{m}$. If the algorithm terminates, it correctly returns such a basis.
@@
Assume that Algorithm 3 terminates with output $\{f_{1},\dots,f_{l}\}$. By construction, these polynomials are generators of $A$ because the set contains the original generators. Further, the $f_{j}$ are $\Delta$/̄homogeneous as the algorithm only adds homogeneous elements. If the algorithm terminates, we must have $J_{i}^{\mathrm{hom}}\subseteq I^{\mathrm{hom}}+\langle t_{i}\rangle$ for all $i\in\{1,\dots,m\}$ for the ‘updated’ ideals corresponding to the $f_{j}$. Hence $\{f_{1},\dots,f_{l}\}$ is a $\Delta$/̄homogeneous MUVAK basis with respect to $v_{1},\dots,v_{m}$ by Theorem 3.13. In particular, $A$ admits such a finite basis.
@@
Conversely, assume that $\{\tilde{f}_{1},\dots,\tilde{f}_{s}\}\subseteq A$ is a finite $\Delta$/̄homogeneous MUVAK basis with respect to $v_{1},\dots,v_{m}$. If each $\tilde{f}_{i}$, $1\leq i\leq s$, is faithfully representable in the elements of the set $\mathcal{B}$ at some point of the algorithm, then $\mathcal{B}$ is a $\Delta$/̄homogeneous MUVAK basis of $A$ with respect to $v_{1},\dots,v_{m}$ by Lemma 3.15 and the algorithm terminates by Theorem 3.13. It hence remains to show that $\tilde{f}_{1},\dots,\tilde{f}_{s}$ are faithfully representable in the elements of the set $\mathcal{B}$ after finitely many iterations.
@@
Let $i\in\{1,\dots,m\}$ such that at the beginning of the $i$/̄th iteration of the for/loop there is $\sigma\in\{1,\dots,s\}$ and $h\in\alpha^{-1}(\tilde{f}_{\sigma})$ with $\operatorname{in}_{i}(h)^{\mathrm{hom}}\in J_{i}^{\mathrm{hom}}\setminus(I^{\mathrm{hom}}+\langle t_{i}\rangle)$. If no such $i$ and $\sigma$ exist, then $\tilde{f}_{1},\dots,\tilde{f}_{s}$ are faithfully representable in the elements of $\mathcal{B}$ by Lemma 3.12. Write =$\mathcal{B}=\{f_{1},\dots,f_{l}\}$ and $J_{i}^{\mathrm{hom}}=\langle h_{1},\dots,h_{r}\rangle$ as in the algorithm. After reordering, we may assume that $h_{1},\dots,h_{r^{\prime}}\notin I^{\mathrm{hom}}+\langle t_{i}\rangle$ and $h_{r^{\prime}+1},\dots,h_{r}\in I^{\mathrm{hom}}+\langle t_{i}\rangle$ for some $r^{\prime}\leq r$. The algorithm now adds $\alpha(h_{j}^{\mathrm{deh}})$ for $j=1,\dots,r^{\prime}$ to the set $\mathcal{B}$. Notice that $h_{j}^{\mathrm{deh}}\in J_{i}$, so $\operatorname{mindeg}_{i}(h_{j}^{\mathrm{deh}})<v_{i}(\alpha(h^{\mathrm{deh}}))$ for $j=1,\dots,r^{\prime}$ by Lemma 2.15. Let $\alpha^{+},I^{+},J_{i}^{+}$ and so on be the updated instances at the end of this iteration of the for/loop. We claim that $h_{1},\dots,h_{r}\in(I^{+})^{\mathrm{hom}}+\langle t_{i}\rangle$. Let $X_{l+1},\dots,X_{l+r^{\prime}}$ be the new variables corresponding to $\alpha(h_{1}^{\mathrm{deh}}),\dots,\alpha(h_{r^{\prime}}^{\mathrm{deh}})$ respectively. Notice that $\alpha^{+}(f^{\prime})=\alpha(f^{\prime})$ for any $f^{\prime}\in K[X_{1},\dots,X_{l}]$ and in particular $I\subseteq I^{+}$ by considering elements of $K[X_{1},\dots,X_{l}]$ as elements of $K[X_{1},\dots,X_{l+r^{\prime}}]$ in the canonical way. For $j\in\{1,\dots,r^{\prime}\}$, we have by construction $h_{j}^{\mathrm{deh}}-X_{l+j}\in I^{+}$, $\deg_{i}(X_{l+j})>\deg_{i}(h_{j})$ and $\deg_{i^{\prime}}(X_{l+j})\geq\deg_{i^{\prime}}(h_{j})$ for all $i^{\prime}\neq i$. Hence $(h_{j}^{\mathrm{deh}}-X_{l+j})^{\mathrm{hom}}=h_{j}-X_{l+j}\vartheta$ with $\vartheta$ a monomial in $t_{1},\dots,t_{m}$ and $t_{i}\mid\vartheta$. So, $h_{j}=(h_{j}^{\mathrm{deh}}-X_{l+j})^{\mathrm{hom}}+X_{l+j}\vartheta\in(I^{+})^{\mathrm{hom}}+\langle t_{i}\rangle$ as claimed. Then also $\operatorname{in}_{i}(h)^{\mathrm{hom}}\in(I^{+})^{\mathrm{hom}}+\langle t_{i}\rangle$. Hence we may find a preimage $h^{\prime}\in\alpha^{-1}(\tilde{f}_{\sigma})$ with $\operatorname{mindeg}_{i}(h^{\prime})>\operatorname{mindeg}_{i}(h)$ and $\operatorname{mindeg}_{i^{\prime}}(h^{\prime})\geq\operatorname{mindeg}_{i^{\prime}}(h)$ for $i^{\prime}\neq i$ by the same argument as in Lemma 3.12. Continuing with the argument as in Lemma 3.12, the algorithm will produce generators in which $\tilde{f}_{1},\dots,\tilde{f}_{s}$ are faithfully representable after finitely many steps. ∎

### v2

@@ 3.3. Computing homogeneous MUVAK bases
Let $h\in\alpha^{-1}(f)$. If $\operatorname{mindeg}_{i}(h)=v_{i}(f)$ for all $i\in\{1,\dots,m\}$, then $f$ is faithfully representable in $f_{1},\dots,f_{k}$ via $h$. So, assume there is $i\in\{1,\dots,m\}$ with $\operatorname{mindeg}_{i}(h)<v_{i}(f)$, hence $\operatorname{in}_{i}(h)^{\mathrm{hom}}\in I^{\mathrm{hom}}+\langle t_{i}\rangle$ by assumption. Then there is $h^{\prime}\in\langle t_{i}\rangle$ with $\operatorname{in}_{i}(h)^{\mathrm{hom}}-h^{\prime}\in I^{\mathrm{hom}}$, so $\operatorname{in}_{i}(h)-h^{\prime\mathrm{deh}}\in I$. We may assume $\deg_{j}(h^{\prime})=\deg_{j}(\operatorname{in}_{i}(h)^{\mathrm{hom}})$ for all $j\in\{1,\dots,m\}$, so $\operatorname{mindeg}_{j}(h^{\prime\mathrm{deh}})\geq\operatorname{mindeg}_{j}(\operatorname{in}_{i}(h))$ and $\operatorname{mindeg}_{i}(h^{\prime\mathrm{deh}})>\deg_{i}(\operatorname{in}_{i}(h))$ as $t_{i}\mid h^{\prime}$. We now replace $h$ by $h-\operatorname{in}_{i}(h)+h^{\prime\mathrm{deh}}$. By continuing to increase the $\deg_{i}$-degree in the described way, we arrive at a polynomial $\tilde{h}\in\alpha^{-1}(f)$ with $\operatorname{mindeg}_{i}(\tilde{h})=v_{i}(f)$ and $\operatorname{mindeg}_{j}(\tilde{h})\geq\operatorname{mindeg}_{j}(h)$ for all $j\neq i$ after finitely many steps. We may now continue this process with another index $i^{\prime}$ for which $\operatorname{mindeg}_{i^{\prime}}(\tilde{h})<v_{i^{\prime}}(f)$ and eventually we obtain a polynomial $h^{\prime}\in R$ with $\alpha(h^{\prime})=f$ and $\operatorname{mindeg}_{j}(h^{\prime})=v_{j}(f)$ for all $j\in\{1,\dots,m\}$, so $f$ is faithfully representable. ∎
@@
Let $f_{1},\dots,f_{k}\in A$ be $\Delta$-homogeneous generators and let $I\coloneqq\ker(\alpha)$ and $J_{i}\coloneqq\ker(\beta_{i})$, $1\leq i\leq m$, with the morphisms $\alpha$ and $\beta_{i}$ as above. The set $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous MUVAK basis of $A$ with respect to $v_{1},\dots,v_{m}$ if and only if we have $J_{i}^{\mathrm{hom}}\subseteq I^{\mathrm{hom}}+\langle t_{i}\rangle$ for all $i\in\{1,\dots,m\}$.
@@
Assume that $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous MUVAK basis of $A$ with respect to $v_{1},\dots,v_{m}$. Fix $i\in\{1,\dots,m\}$ and let $h\in J_{i}$. As $J_{i}$ is a $\Delta$-homogeneous ideal, we may assume that $h$ is $\Delta$-homogeneous. Set $f\coloneqq h(f_{1},\dots,f_{k})$. By assumption, the polynomial $f$ is faithfully representable, so by Lemma 3.14, we have $h^{\mathrm{hom}}\in I^{\mathrm{hom}}+\langle\vartheta\rangle$ with $\vartheta=\prod_{j=1}^{m}t_{j}^{v_{j}(f)-\deg_{j}(h^{\mathrm{hom}})}$. By Lemma 2.17, we have $\operatorname{mindeg}_{i}(h)<v_{i}(f)$, so $t_{i}\mid\vartheta$. Hence $h^{\mathrm{hom}}\in I^{\mathrm{hom}}+\langle t_{i}\rangle$.
Assume now $J_{i}^{\mathrm{hom}}\subseteq I^{\mathrm{hom}}+\langle t_{i}\rangle$ for all $i\in\{1,\dots,m\}$. Let $f\in A$ be $\Delta$-homogeneous and pick $h\in\alpha^{-1}(f)$. If $\operatorname{mindeg}_{i}(h)\geq v_{i}(f)$ for all $i\in\{1,\dots,m\}$, then $f$ is faithfully representable, so assume that there is $i\in\{1,\dots,m\}$ with $\operatorname{mindeg}_{i}(h)<v_{i}(f)$. Then $\operatorname{in}_{i}(h)\in J_{i}$ by Lemma 2.17, so $\operatorname{in}_{i}(h)^{\mathrm{hom}}\in J_{i}^{\mathrm{hom}}\subseteq I^{\mathrm{hom}}+\langle t_{i}\rangle$. Hence $f$ is faithfully representable in $f_{1},\dots,f_{k}$ with respect to $v_{1},\dots,v_{m}$ by Lemma 3.16. ∎
Example 3.18.
This example is part of the series also containing 3.3, 3.15 and 3.22. Computing the relations of the initial forms gives the ideals
$J_{1}=\langle X_{3}-X_{4},X_{1}^{2}-X_{2}X_{4}\rangle\text{ and }J_{2}=\langle X_{1}-X_{2},X_{3}^{2}-X_{2}X_{4}\rangle,$
where $J_{1}$ corresponds to $v_{x}$ and $J_{2}$ to $v_{y}$. Homogenizing yields
$J_{1}^{\mathrm{hom}}=\langle X_{3}-X_{4}t_{2},X_{1}^{2}-X_{2}X_{4}t_{2}^{2}\rangle\text{ and }J_{2}^{\mathrm{hom}}=\langle X_{1}-X_{2}t_{1},X_{3}^{2}-X_{2}X_{4}t_{1}^{2}\rangle.$
One checks that $J_{i}^{\mathrm{hom}}\not\subseteq I^{\mathrm{hom}}+\langle t_{i}\rangle$ for $i=1,2$.
@@
Input : $\Delta$-homogeneous generators $f_{1},\dots,f_{k}\in A\leq B$, valuations $v_{1},\dots,v_{m}:B\setminus\{0\}\to\mathbb{Z}$
Output : A $\Delta$-homogeneous MUVAK basis of $A$ with respect to $v_{1},\dots,v_{m}$
1
21exInitialize $\mathcal{G}\coloneqq\{f_{1},\dots,f_{k}\}$
@@ 3 Compute the ideals $I^{\mathrm{hom}}$ and $J_{i}^{\mathrm{hom}}$ for all $i\in\{1,\dots,m\}$
6     Write $J_{i}^{\mathrm{hom}}=\langle h_{1},\dots,h_{r}\rangle$ and $\mathcal{G}=\{f_{1},\dots,f_{l}\}$
7     $\mathcal{G}\coloneqq\mathcal{G}\cup\{h_{j}^{\mathrm{deh}}(f_{1},\dots,f_{l})\mid h_{j}\notin I^{\mathrm{hom}}+\langle t_{i}\rangle,\penalty 10000\ j=1,\dots,r\}$
@@ 8     Update $I^{\mathrm{hom}}$ and $J_{i}^{\mathrm{hom}}$
We give an efficient method to compute the homogenization of an ideal with respect to the in general only non-negative weights $v_{i}(f_{j})$ in Appendix A.
@@
Let $\{\tilde{f}_{1},\dots,\tilde{f}_{s}\}\subseteq A$ be a $\Delta$-homogeneous MUVAK basis of $A$ with respect to $v_{1},\dots,v_{m}$ and let $f_{1},\dots,f_{k}\in A$ be $\Delta$-homogeneous generators. If $\tilde{f}_{j}$ is faithfully representable in $f_{1},\dots,f_{k}$ for all $j\in\{1,\dots,s\}$, then $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous MUVAK basis for $A$ with respect to $v_{1},\dots,v_{m}$.
@@
Let $f\in A$ be $\Delta$-homogeneous. Then $f$ is faithfully representable in $\tilde{f}_{1},\dots,\tilde{f}_{s}$, so there is $h\in K[Y_{1},\dots,Y_{s}]$ with $h(\tilde{f}_{1},\dots,\tilde{f}_{s})=f$ and $\operatorname{mindeg}_{i}(h)=v_{i}(f)$ for all $i$. Put $\tilde{h}\coloneqq h(h_{1},\dots,h_{s})\in K[X_{1},\dots,X_{k}]$, so $\tilde{h}(f_{1},\dots,f_{k})=f$. By definition, we have $\deg_{i}(Y_{j})=v_{i}(\tilde{f}_{j})$ for all $i=1,\dots,m$ and $j=1,\dots,s$. Hence $\operatorname{mindeg}_{i}(\tilde{h})\geq\operatorname{mindeg}_{i}(h)$. We conclude $\operatorname{mindeg}_{i}(\tilde{h})\geq v_{i}(f)$ for all $i$, so $f$ is faithfully representable in $f_{1},\dots,f_{k}$ as required. ∎
@@
Algorithm 3 terminates after finitely many steps if and only if there exists a finite $\Delta$-homogeneous MUVAK basis of $A$ with respect to $v_{1},\dots,v_{m}$. If the algorithm terminates, it correctly returns such a basis.
@@
Assume that Algorithm 3 terminates with output $\{f_{1},\dots,f_{l}\}$. By construction, these polynomials are generators of $A$ because the set contains the original generators. Further, the $f_{j}$ are $\Delta$-homogeneous as the algorithm only adds homogeneous elements. If the algorithm terminates, we must have $J_{i}^{\mathrm{hom}}\subseteq I^{\mathrm{hom}}+\langle t_{i}\rangle$ for all $i\in\{1,\dots,m\}$ for the ‘updated’ ideals corresponding to the $f_{j}$. Hence $\{f_{1},\dots,f_{l}\}$ is a $\Delta$-homogeneous MUVAK basis with respect to $v_{1},\dots,v_{m}$ by Theorem 3.17. In particular, $A$ admits such a finite basis.
@@
Conversely, assume that $\{\tilde{f}_{1},\dots,\tilde{f}_{s}\}\subseteq A$ is a finite $\Delta$-homogeneous MUVAK basis with respect to $v_{1},\dots,v_{m}$. If each $\tilde{f}_{i}$, $1\leq i\leq s$, is faithfully representable in the elements of the set $\mathcal{G}$ at some point of the algorithm, then $\mathcal{G}$ is a $\Delta$-homogeneous MUVAK basis of $A$ with respect to $v_{1},\dots,v_{m}$ by Lemma 3.20 and the algorithm terminates by Theorem 3.17. It hence remains to show that $\tilde{f}_{1},\dots,\tilde{f}_{s}$ are faithfully representable in the elements of the set $\mathcal{G}$ after finitely many iterations.
@@
Let $i\in\{1,\dots,m\}$ such that at the beginning of the $i$-th iteration of the for-loop there is $\sigma\in\{1,\dots,s\}$ and $h\in\alpha^{-1}(\tilde{f}_{\sigma})$ with $\operatorname{in}_{i}(h)^{\mathrm{hom}}\in J_{i}^{\mathrm{hom}}\setminus(I^{\mathrm{hom}}+\langle t_{i}\rangle)$. If no such $i$ and $\sigma$ exist, then $\tilde{f}_{1},\dots,\tilde{f}_{s}$ are faithfully representable in the elements of $\mathcal{G}$ by Lemma 3.16. Write $\mathcal{G}=\{f_{1},\dots,f_{l}\}$ and $J_{i}^{\mathrm{hom}}=\langle h_{1},\dots,h_{r}\rangle$ as in the algorithm. After reordering, we may assume that $h_{1},\dots,h_{r^{\prime}}\notin I^{\mathrm{hom}}+\langle t_{i}\rangle$ and $h_{r^{\prime}+1},\dots,h_{r}\in I^{\mathrm{hom}}+\langle t_{i}\rangle$ for some $r^{\prime}\leq r$. The algorithm now adds $\alpha(h_{j}^{\mathrm{deh}})$ for $j=1,\dots,r^{\prime}$ to the set $\mathcal{G}$. Notice that $h_{j}^{\mathrm{deh}}\in J_{i}$, so $\operatorname{mindeg}_{i}(h_{j}^{\mathrm{deh}})<v_{i}(\alpha(h^{\mathrm{deh}}))$ for $j=1,\dots,r^{\prime}$ by Lemma 2.17. Let $\alpha^{+},I^{+},J_{i}^{+}$ and so on be the updated instances at the end of this iteration of the for-loop. We claim that $h_{1},\dots,h_{r}\in(I^{+})^{\mathrm{hom}}+\langle t_{i}\rangle$. Let $X_{l+1},\dots,X_{l+r^{\prime}}$ be the new variables corresponding to $\alpha(h_{1}^{\mathrm{deh}}),\dots,\alpha(h_{r^{\prime}}^{\mathrm{deh}})$ respectively. Notice that $\alpha^{+}(f^{\prime})=\alpha(f^{\prime})$ for any $f^{\prime}\in K[X_{1},\dots,X_{l}]$ and in particular $I\subseteq I^{+}$ by considering elements of $K[X_{1},\dots,X_{l}]$ as elements of $K[X_{1},\dots,X_{l+r^{\prime}}]$ in the canonical way. For $j\in\{1,\dots,r^{\prime}\}$, we have by construction $h_{j}^{\mathrm{deh}}-X_{l+j}\in I^{+}$, $\deg_{i}(X_{l+j})>\deg_{i}(h_{j})$ and $\deg_{i^{\prime}}(X_{l+j})\geq\deg_{i^{\prime}}(h_{j})$ for all $i^{\prime}\neq i$. Hence $(h_{j}^{\mathrm{deh}}-X_{l+j})^{\mathrm{hom}}=h_{j}-X_{l+j}\vartheta$ with $\vartheta$ a monomial in $t_{1},\dots,t_{m}$ and $t_{i}\mid\vartheta$. So, $h_{j}=(h_{j}^{\mathrm{deh}}-X_{l+j})^{\mathrm{hom}}+X_{l+j}\vartheta\in(I^{+})^{\mathrm{hom}}+\langle t_{i}\rangle$ as claimed. Then also $\operatorname{in}_{i}(h)^{\mathrm{hom}}\in(I^{+})^{\mathrm{hom}}+\langle t_{i}\rangle$. Hence we may find a preimage $h^{\prime}\in\alpha^{-1}(\tilde{f}_{\sigma})$ with $\operatorname{mindeg}_{i}(h^{\prime})>\operatorname{mindeg}_{i}(h)$ and $\operatorname{mindeg}_{i^{\prime}}(h^{\prime})\geq\operatorname{mindeg}_{i^{\prime}}(h)$ for $i^{\prime}\neq i$ by the same argument as in Lemma 3.16. Continuing with the argument as in Lemma 3.16, the algorithm will produce generators in which $\tilde{f}_{1},\dots,\tilde{f}_{s}$ are faithfully representable after finitely many steps. ∎
Example 3.22.
This example is part of the series also containing 3.3, 3.15 and 3.18. We use Algorithm 3 to compute a MUVAK basis of $A$. Let $\mathcal{G}=\{f_{1},\dots,f_{4}\}$. With Example 3.18, we have the polynomials $h_{1}=X_{3}-X_{4}t_{2},h_{2}=X_{1}^{2}-X_{2}X_{4}t_{2}^{2}\notin I^{\mathrm{hom}}+\langle t_{1}\rangle$ and $h_{3}=X_{1}-X_{2}t_{1}\notin I^{\mathrm{hom}}+\langle t_{2}\rangle$ (notice that the second generator of $J_{2}^{\mathrm{hom}}$ is indeed an element of $I^{\mathrm{hom}}+\langle t_{2}\rangle$). These give the polynomials
$h_{1}^{\mathrm{deh}}(f_{1},\dots,f_{4})=h_{3}^{\mathrm{deh}}(f_{1},\dots,f_{4})=xy\text{ and }h_{2}^{\mathrm{deh}}(f_{1},\dots,f_{4})=x^{4}+2x^{3}y.$
For simplicity, we only add $f_{5}\coloneqq xy$ to the set of generators. Computing the new relations gives the ideals
$I^{\mathrm{hom}}$ $=\langle X_{3}-X_{4}t_{2}-X_{5}t_{1},X_{1}-X_{2}t_{1}-X_{5}t_{2},X_{2}X_{4}-X_{5}^{2}\rangle,$
$J_{1}^{\mathrm{hom}}$ $=\langle X_{3}-X_{4}t_{2},X_{1}-X_{5}t_{2},X_{2}X_{4}-X_{5}^{2}\rangle,$
$J_{2}^{\mathrm{hom}}$ $=\langle X_{3}-X_{5}t_{1},X_{1}-X_{2}t_{1},X_{2}X_{4}-X_{5}^{2}\rangle,$
where the homogenization is now with respect to the matrix
$W=\begin{pmatrix}1&2&0&0&1\\ 0&0&1&2&1\end{pmatrix}$.
We see that now $J_{i}^{\mathrm{hom}}\subseteq I^{\mathrm{hom}}+\langle t_{i}\rangle$ and the algorithm terminates with the MUVAK basis $\{f_{1},\dots,f_{5}\}$ for $A$.